OpenStudy (anonymous):

Simplify each expression: 0.6(9.7-1.5v)+4(1+3.4v) i need the steps. I know the answers 9.82+12.7v

OpenStudy (anonymous):

5.82 - 0.9v + 4 + 13.6v =9.82 + 12.7v

OpenStudy (anonymous):

Multiply the constants with the terms in parenthesis and then add the like terms

OpenStudy (anonymous):

So that would be \[(0.6 \times 9.7)+(0.6\times -1.5v)+(4\times1)+(4\times 3.4v)\]